# Classical billiards can compute

### v1

@@ 1. Introduction
Given the richness of billiard dynamics, a natural question arises: can billiards realize the full complexity of computation? Since the pioneering work of Moore [24, 25], it has become clear that many continuous dynamical systems can simulate algorithms and thereby exhibit undecidability phenomena rooted in Turing’s Halting problem [30]. Universal computation has since been embedded in a wide range of settings, including polynomial differential equations, mechanical systems with tailored potentials, cellular automata and hydrodynamic flows [22, 18, 8, 31, 32, 29, 9, 12, 13, 11, 10, 17]. A unifying framework behind several of these results is Topological Kleene Field Theory, which encodes computable functions into flows on manifolds with boundary so that dynamical questions translate into questions of computation [16].
@@
Acknowledgements: We thank FIM and ETH Zürich for providing an ideal environment for this collaboration. We are also grateful to Stephen Wolfram for his visit in October 2025, which inspired several related questions that ultimately led to this work. We additionally thank Ángel González-Prieto for insightful discussions that shaped our encoding of the computational states of a Turing machine into a one-dimensional interval.
@@ 3.1. Cantor-based encodings
We would also like to add somehow the information about the position of the head into this encoding. In order to do so, we will embed a smaller copy of $I$ into itself for each $k\in\mathbb{Z}$ via an affine map $\tau_{k}:I\hookrightarrow I$. By having the length of $I_{k}:=\tau_{k}(I)$ decrease fast enough as $|k|$ goes to infinity, we can construct these embeddings so that the intervals $I_{k}$ are pairwise disjoint. If we also assume that $I_{k}$ is to the left of $I_{k^{\prime}}$ if $k<k^{\prime}$ and we have symmetry with respect to the center of the interval, in the sense that $I_{k}=-I_{-k}$, the result we obtain looks like the one in Figure 1. Using these intervals, we can encode the tape $t$ when the head state is at position $k$ as the point $x_{t,k}:=\tau_{k}(x_{t})\in[-1,1]$.
@@ 3.3. Proof of the main theorem
Notice that, since there are finite gaps between the intervals where the computation states are, these straight segments can be joined smoothly, producing a differentiable curve. Moreover, although the number of alternating intervals grows with $|k|$, their lengths decrease more rapidly (otherwise they would not fit inside the compact interval $I$). As a result, if we observe the curve over $I_{k}$ as $|k|\to\infty$, it resembles the graph of $f(x)\sin(1/g(x))$ as $x\to 0$, where $f(x),g(x)\to 0$ with $f(x)\ll g(x)$ as $x\to 0$. This makes it a perfectly valid billiard wall, as illustrated in Figure 6.

### v2

@@ 1. Introduction
Given the richness of billiard dynamics, a natural question arises: can billiards realize the full complexity of computation? Since the pioneering work of Moore [25, 26], it has become clear that many continuous dynamical systems can simulate algorithms and thereby exhibit undecidability phenomena rooted in Turing’s Halting problem [31]. At a more foundational level, the question of whether physical systems can genuinely realize such Turing-universal behavior has been central since the very origins of computation, notably in the seminal works of Feynman [15] and Wolfram [33]. Universal computation has since been embedded in a wide range of settings, including polynomial differential equations, mechanical systems with tailored potentials, cellular automata and hydrodynamic flows [23, 19, 8, 32, 34, 30, 9, 12, 13, 11, 10, 18]. A unifying framework behind several of these results is Topological Kleene Field Theory, which encodes computable functions into flows on manifolds with boundary so that dynamical questions translate into questions of computation [17].
@@
Acknowledgements: We thank FIM and ETH Zürich for providing an ideal environment for this collaboration. We are also grateful to Stephen Wolfram for his visit in October 2025, which inspired several related questions that ultimately led to this work. We further thank Ángel González-Prieto for insightful discussions that shaped our encoding of the computational states of a Turing machine into a one-dimensional interval. We also thank John Baez and Joel David Hamkins for their interest.
@@ 3.1. Cantor-based encodings
We would also like to add somehow the information about the position of the head into this encoding. In order to do so, we will embed a smaller copy of $I$ into itself for each $k\in\mathbb{Z}$ via an affine map $\tau_{k}:I\hookrightarrow I$. By having the length of $I_{k}:=\tau_{k}(I)$ decrease fast enough as $|k|$ goes to infinity, we can construct these embeddings so that the intervals $I_{k}$ are pairwise disjoint. If we also assume that $I_{k}$ is to the left of $I_{k^{\prime}}$ if $k<k^{\prime}$ and we have symmetry with respect to the center of the interval, the result we obtain looks like the one in Figure 1. Using these intervals, we can encode the tape $t$ when the head state is at position $k$ as the point $x_{t,k}:=\tau_{k}(x_{t})\in[0,1]$.
@@ 3.3. Proof of the main theorem
Notice that, since there are finite gaps between the intervals where the computation states are, these straight segments can be joined smoothly, producing a differentiable curve.
For a more thorough explanation of how this curve is constructed, refer to the Appendix.
@@ 5. Conclusions
Appendix
The purpose of this appendix is to provide the missing details in the proof of Lemma 2 from the main text, thus completing the argument that the read–write operation can be realized by means of billiard dynamics.
Recall first that, according to our definition of a Turing machine, applying the read–write operation to a computation state $(t,k)$ requires inspecting and possibly modifying the tape symbol $t_{k}$. Under our Cantor set–type encoding, this amounts to detecting and, if necessary, changing a specific digit in the ternary expansion associated with the tape before applying the embedding $\tau_{k}$. For this reason, we argued that, inside the interval $I_{k}$ corresponding to the head position $k$, the admissible computation states decompose into two alternating families of subintervals, depending on whether the symbol at position $k$ is 0 or $1$. We now make this decomposition explicit.
Fix $k\in\mathbb{Z}$ and, for simplicity, assume $k\geq 0$ (the case $k<0$ is completely analogous). We begin by considering the encoding of the tape $t$ alone as a point $x_{t}\in I$, prior to applying the embedding $\tau_{k}$. Our goal is to describe two families of Cantor set blocks corresponding to the conditions $t_{k}=0$ and $t_{k}=1$. Recall that, in our encoding, the symbol $t_{k}$ is stored in the $(2k+1)$-th digit of the ternary expansion of $x_{t}$. To simplify notation, we introduce
$x_{\epsilon_{1}\ldots\epsilon_{l}}:=2\sum_{n=1}^{l}\epsilon_{n}\,3^{-n},$
where each $\epsilon_{n}\in\{0,1\}$.
With this notation, the Cantor set blocks corresponding to tapes with $t_{k}=0$ are given by
$I_{\epsilon_{1}\ldots\epsilon_{2k}0}=\bigl[x_{\epsilon_{1}\ldots\epsilon_{2k}0},\,x_{\epsilon_{1}\ldots\epsilon_{2k}0}+3^{-(2k+1)}\bigr],\qquad\epsilon_{n}\in\{0,1\},\ n=1,\ldots,2k,$
while those corresponding to tapes with $t_{k}=1$ are
$I_{\epsilon_{1}\ldots\epsilon_{2k}1}=\bigl[x_{\epsilon_{1}\ldots\epsilon_{2k}1},\,x_{\epsilon_{1}\ldots\epsilon_{2k}1}+3^{-(2k+1)}\bigr],\qquad\epsilon_{n}\in\{0,1\},\ n=1,\ldots,2k.$
We now apply the embedding $\tau_{k}$, using its explicit expression from the proof of Lemma 1. Define
$x_{\epsilon_{1}\ldots\epsilon_{l}}^{k}:=\tau_{k}(x_{\epsilon_{1}\ldots\epsilon_{l}})=1+3^{-(1+k)}\bigl(-2+x_{\epsilon_{1}\ldots\epsilon_{l}}\bigr).$
Then the Cantor set blocks inside $I_{k}$ corresponding to the condition $t_{k}=s$, with $s\in\{0,1\}$, are
$I_{\epsilon_{1}\ldots\epsilon_{2k}s}^{k}:=\tau_{k}\bigl(I_{\epsilon_{1}\ldots\epsilon_{2k}s}\bigr)=\bigl[x_{\epsilon_{1}\ldots\epsilon_{2k}s}^{k},\,x_{\epsilon_{1}\ldots\epsilon_{2k}s}^{k}+3^{-(3k+2)}\bigr],$
again with $\epsilon_{n}\in\{0,1\}$ for $n=1,\ldots,2k$. There are $2^{2k}$ such intervals, each of length $3^{-(3k+2)}$.
We now describe the billiard walls implementing the read–write operation. We assume that incoming billiard trajectories move upward, with positive velocity in the $y$-direction, and have horizontal coordinate $x\in[0,1]$, where the computation state $(t,k)$ is encoded as $x_{t,k}$. We first treat the case in which no symbol change occurs, so that the goal is simply to distinguish between $t_{k}=0$ and $t_{k}=1$.
After interacting with the billiard walls, we want trajectories corresponding to $t_{k}=0$ to continue vertically with horizontal coordinate $x_{t,k}-2$, while those corresponding to $t_{k}=1$ should continue vertically with horizontal coordinate $x_{t,k}+2$. To achieve this separation, trajectories first collide with straight walls placed above the intervals $I_{\epsilon_{1}\ldots\epsilon_{2k}s}^{k}$, with slope 1-1 for $s=0$ and slope +1+1 for $s=1$.
For $t_{k}=0$, we define the walls
$W_{\epsilon_{1}\ldots\epsilon_{2k}0}^{k}(x)=2+x_{\epsilon_{1}\ldots\epsilon_{2k}0}^{k}+\frac{3^{-(3k+2)}}{2}-x,\qquad x\in I_{\epsilon_{1}\ldots\epsilon_{2k}0}^{k}.$
Each such wall is a line segment of slope 1-1, centered horizontally above the corresponding interval $I_{\epsilon_{1}\ldots\epsilon_{2k}0}^{k}$ and located at height $y=2$.
Similarly, for $t_{k}=1$, we take walls of slope +1+1 given by
$W_{\epsilon_{1}\ldots\epsilon_{2k}1}^{k}(x)=2-x_{\epsilon_{1}\ldots\epsilon_{2k}1}^{k}-\frac{3^{-(3k+2)}}{2}+x,\qquad x\in I_{\epsilon_{1}\ldots\epsilon_{2k}1}^{k}.$
These walls deflect trajectories corresponding to $t_{k}=0$ to the left and those corresponding to $t_{k}=1$ to the right.
Once the trajectories have been separated, we must redirect them so that they again travel vertically. For instance, in the case $t_{k}=0$, this is achieved by introducing a second family of walls
$\widetilde{W}_{\epsilon_{1}\ldots\epsilon_{2k}0}^{k}(x)=x_{\epsilon_{1}\ldots\epsilon_{2k}0}^{k}+\frac{3^{-(3k+2)}}{2}-(x+2),$
defined for
$x\in\bigl[x_{\epsilon_{1}\ldots\epsilon_{2k}0}^{k}-2,\,x_{\epsilon_{1}\ldots\epsilon_{2k}0}^{k}-2+3^{-(3k+2)}\bigr].$
A trajectory starting at $x_{t,k}$ with $t_{k}=0$ first reflects off the wall $W_{\epsilon_{1}\ldots\epsilon_{2k}0}^{k}$, then travels along a line of slope 1-1 until it reaches $\widetilde{W}_{\epsilon_{1}\ldots\epsilon_{2k}0}^{k}$. Since the two walls are parallel, the outgoing trajectory is vertical, with horizontal coordinate $x_{t,k}-2$, as desired (see Figure 6). A similar construction can be used to have the outgoing trajectory corresponding to $t_{k}=1$ travel vertically with horizontal coordinate $x_{t,k}+2$.
Figure 6. Billiard dynamics implementing the read–write operation with no symbol change. The blue and red rays correspond to reading a 0 or a 1, respectively.
We now turn to the case in which a symbol change is required. For simplicity, suppose that we want to modify the tape symbol at position $k$ from $t_{k}=0$ to $t^{\prime}_{k}=1$. In terms of our encoding, this means that an incoming vertical trajectory with horizontal coordinate $x_{t,k}$ should be transformed, after interacting with the read–write construction, into a vertical trajectory with horizontal coordinate
$x_{t,k}+2\cdot 3^{-(3k+2)}-2.$
Compared with the pure read case, this requires a slight horizontal displacement of the trajectory. Therefore, the separating wall can no longer have slope exactly 1-1. Instead, we choose the angle $\alpha_{k}$ with the horizontal so that
$\tan\alpha_{k}=\frac{1}{1+3^{-(3k+2)}}.$
This deviation from slope 1-1 is very small, especially for large values of $k$, and in the limit $k\to\infty$ we recover the original configuration. We can observe this in Figure 7, where the small displacement is displayed compared to the original no-rewrite wall.
Figure 7. Billiard dynamics implementing the read–write operation with symbol change compared to the no-change situation.
We must now verify that all these trajectories travel from the walls $W_{\epsilon_{1}\ldots\epsilon_{2k}s}^{k}$ to the corresponding $\widetilde{W}_{\epsilon_{1}\ldots\epsilon_{2k}s}^{k}$ without colliding with any other wall. Since all walls lie at the same height, this reduces to checking that, after bouncing off $W_{\epsilon_{1}\ldots\epsilon_{2k}s}^{k}$, the trajectory descends sufficiently fast so as not to intersect any other wall $W_{\epsilon^{\prime}_{1}\ldots\epsilon^{\prime}_{2k^{\prime}}s^{\prime}}^{k^{\prime}}$.
Observe first that trajectories corresponding to a symbol change have a steeper downward slope than those in the no-rewrite case. As a result, avoiding collisions is strictly easier in the rewriting case. It therefore suffices to carry out the check for the read–only configuration, where the walls have slopes $\pm 1$.
Consider a trajectory that bounces off a wall $W_{\epsilon_{1}\ldots\epsilon_{2k}0}^{k}$. Until it reaches $\widetilde{W}_{\epsilon_{1}\ldots\epsilon_{2k}0}^{k}$, this trajectory remains parallel to all walls of the form $W_{\epsilon^{\prime}_{1}\ldots\epsilon^{\prime}_{2k^{\prime}}1}^{k^{\prime}}$ (or their slope is larger when the symbol $1$ is modified by the transition). Since such walls are separated by a horizontal distance of at least $2\cdot 3^{-(3k+2)}$, there is no risk of collision with them. We therefore only need to check that the trajectory does not hit any wall $W_{\epsilon^{\prime}_{1}\ldots\epsilon^{\prime}_{2k^{\prime}}0}^{k^{\prime}}$ lying to its left.
It is enough to consider the extremal situation. Namely, we check that the trajectory bouncing off the left endpoint of $W_{\epsilon_{1}\ldots\epsilon_{2k}0}^{k}$, located at
$\bigl(x_{\epsilon_{1}\ldots\epsilon_{2k}0}^{k},\,2+3^{-(3k+2)}/2\bigr),$
does not intersect the right endpoint of any wall $W_{\epsilon^{\prime}_{1}\ldots\epsilon^{\prime}_{2k^{\prime}}0}^{k^{\prime}}$, located at
$\bigl(x_{\epsilon^{\prime}_{1}\ldots\epsilon^{\prime}_{2k^{\prime}}0}^{k^{\prime}}+3^{-(3k^{\prime}+2)},\,2-3^{-(3k^{\prime}+2)}/2\bigr).$
This amounts to verifying the inequality
$x_{\epsilon_{1}\ldots\epsilon_{2k}0}^{k}-\bigl(x_{\epsilon^{\prime}_{1}\ldots\epsilon^{\prime}_{2k^{\prime}}0}^{k^{\prime}}+3^{-(3k^{\prime}+2)}\bigr)>\frac{3^{-(3k+2)}}{2}+\frac{3^{-(3k^{\prime}+2)}}{2}.$
We first consider the case $k=k^{\prime}$. By construction, if $x_{\epsilon_{1}\ldots\epsilon_{2k}0}^{k}\neq x_{\epsilon^{\prime}_{1}\ldots\epsilon^{\prime}_{2k}0}^{k}$, then
$x_{\epsilon_{1}\ldots\epsilon_{2k}0}^{k}-x_{\epsilon^{\prime}_{1}\ldots\epsilon^{\prime}_{2k}0}^{k}\geq 2\cdot 3^{-(3k+1)}.$
It follows that
$x_{\epsilon_{1}\ldots\epsilon_{2k}0}^{k}-\bigl(x_{\epsilon^{\prime}_{1}\ldots\epsilon^{\prime}_{2k}0}^{k}+3^{-(3k+2)}\bigr)\geq 3^{-(3k+2)}(6-1)>3^{-(3k+2)},$
which is more than enough to guarantee the desired separation.
We now turn to the case $k\neq k^{\prime}$. In this situation, $x_{\epsilon_{1}\ldots\epsilon_{2k}0}^{k}\in I_{k}$ and $x_{\epsilon^{\prime}_{1}\ldots\epsilon^{\prime}_{2k^{\prime}}0}^{k^{\prime}}\in I_{k^{\prime}}$. Since the distance between $I_{k^{\prime}}$ and $I_{k^{\prime}+1}$ is $3^{-(k^{\prime}+2)}$, the distance between $I_{k^{\prime}}$ and $I_{k}$ is at least this large. Consequently,
$x_{\epsilon_{1}\ldots\epsilon_{2k}0}^{k}-\bigl(x_{\epsilon^{\prime}_{1}\ldots\epsilon^{\prime}_{2k^{\prime}}0}^{k^{\prime}}+3^{-(3k^{\prime}+2)}\bigr)\geq 3^{-(k^{\prime}+2)}\geq 3^{-(3k^{\prime}+2)}>\frac{3^{-(3k+2)}}{2}+\frac{3^{-(3k^{\prime}+2)}}{2},$
where in the last inequality we used the fact that $k>k^{\prime}$.
We conclude that, after bouncing off $W_{\epsilon_{1}\ldots\epsilon_{2k}0}^{k}$, the trajectory does not intersect any other wall $W_{\epsilon^{\prime}_{1}\ldots\epsilon^{\prime}_{2k^{\prime}}s^{\prime}}^{k^{\prime}}$. By symmetry, the same argument applies to walls of the form $W_{\epsilon_{1}\ldots\epsilon_{2k}1}^{k}$ and to possible collisions with $\widetilde{W}_{\epsilon_{1}\ldots\epsilon_{2k}s}^{k}$, as well as to the case $k<0$.
Finally, we note that the estimates used above to rule out unwanted collisions are far from sharp. In particular, they leave plenty of room to connect the straight wall segments by means of smooth bump functions, allowing us to construct a completely smooth billiard wall with a zig–zag profile.
Each straight segment in this zig–zag has vertical size on the order of $3^{-(3k+2)}$, and the total number of such segments is $2^{2k}$. As $|k|\to\infty$, the amplitude of the oscillations therefore tends to zero, while their support occupies an increasingly small portion of the interval $I_{k}$. Indeed, the total length of the Cantor blocks of order $2k+1$ (or $2|k|$) vanishes in this limit. Therefore, near the ends of the interval the zig–zag wall converges to a straight line, as shown in Figure 8.